# Exploring Empty Spaces: Human-in-the-Loop Data Augmentation

### v1

@@ 1. Introduction
In machine learning (ML), data plays a key role in driving model behavior. Even the most sophisticated, specialized model architectures may underperform when training data is limited or if data instances are low quality and noisy (Xie et al., 2020; Shorten and Khoshgoftaar, 2019; Sambasivan et al., 2021; Lara and Tiwari, 2022). As such, data augmentation — the process of creating new data samples to add to an existing dataset — is a common and important practice across many ML applications (e.g., model training (Rebuffi et al., 2021; Khosla and Saini, 2020) and fine-tuning (Feng et al., 2020; Shi et al., 2022)). In particular, data augmentation plays a critical role in generating novel use cases for model evaluation, as in the case of perturbation analyses (Shorten and Khoshgoftaar, 2019; Xie et al., 2020; Sui et al., 2024), fairness testing (Navarro et al., 2024; Zhang and Sang, 2020), and red teaming attacks (Feffer et al., 2024; Ganguli et al., 2022). Across many of these scenarios, a shared goal is to improve dataset diversity to make ML models more robust and safe (Chao et al., 2023; Shi et al., 2022; Rebuffi et al., 2021), which is especially important given the growing complexity and applications of these models.
@@
The ease and efficacy of data augmentation is also highly dependent on data modality. For instance, with structured data like tables, it is relatively straightforward to determine where augmentation might be needed by looking at the distribution of features (i.e., categories) and adding new rows. Tabular data can also be augmented by adding entirely new features to the table (i.e., columns). On the other hand, it is much harder to augment unstructured data, like images or text (Figure 1). Such modalities do not have the same kind of identifiable “features” to augment along. Furthermore, for text data, it is unclear how to perform “distortions” to natural language without altering its semantics or context (Ding et al., 2024; Qu et al., 2021), whereas for images, it is easier to apply rotation, cropping, and brightness modifications to augment data. With unstructured text, it is also less clear what kinds of diversity to strive for when performing augmentation, as there are many axes to consider, including topical, syntactic, and lexical diversity (Chen et al., 2020; Reif et al., 2023).
@@
With these augmentation methods, we designed and developed Amplio, an interactive data augmentation tool for unstructured text datasets. Amplio is designed to help ML practitioners systematically navigate “unknown unknowns” and diversify their datasets by finding relevant empty data spaces (i.e., parts of the desired dataset distribution with few or no data points) to explore (Figure 1). Our tool works to support augmentation processes by visualizing sentence embeddings and providing an interface for users to fill in the empty spaces using our three techniques. To assess these augmentation approaches, we conducted a user study designed around red teaming LLMs, a common and important real-world data augmentation task for model evaluation, as identified by our formative study. We recruited 18 professional red teamers to augment a harmful LLM prompts dataset (Bai et al., 2022) using Amplio, finding that our augmentation methods and visualizations were effective in generating relevant, diverse data. Participants also found unique use cases for each augmentation technique, suggesting additional design opportunities for interactive data augmentation tools.
@@
A suite of three human-in-the-loop data augmentation techniques—Augment with Concepts, Augment by Interpolation, and Augment with LLM—to help users find diverse and relevant “empty spaces” to explore when augmenting unstructured text.
The design and implementation of Amplio, an interactive visualization tool that applies our three techniques to help users augment datasets in a controllable and interpretable way.
@@ 2. Related Work
Our system, Amplio, aims to provide a middle ground between freeform and structured text augmentation approaches.
Visualization plays a key role in understanding and evaluating ML models (Brath et al., 2023; Beauxis-Aussalet et al., 2021). Many existing tools focus on exploring data distributions and quality (Research, 2021a, 2017; Face, 2021; IBM, 2021; Assogba et al., 2023; Giesen et al., 2017), understanding data iteration (Hohman et al., 2020), or assessing factors like model fairness (Ahn and Lin, 2019; Bird et al., 2020; Xie et al., 2020; Wexler et al., 2019) and interpretability (Hohman et al., 2019; Tenney et al., 2020; Wang et al., 2020b; Strobelt et al., 2018). There is also interest in designing visualization techniques to aid developers in creating safer, trustworthy, and responsible models (Reif et al., 2023; Templeton, 2024; Beauxis-Aussalet et al., 2021; Feng et al., 2024; Wang et al., 2024).
In contrast, our work targets data augmentation processes, using human-in-the-loop techniques to improve data diversity for unstructured text. Specifically, we focus on synthetic text augmentation.
@@ 2.1. Synthetic Text Data Augmentation
Recently, language models have emerged as an alternative mechanism for data space augmentations (Guo and Chen, 2024; Wagner et al., 2024; Patel et al., 2024; Li et al., 2023; Peng et al., 2023). For example, LLMs have been used to generate datasets of counterfactual sentences (Wu et al., 2021) and adversarial prompts (Samvelyan et al., 2024). While these approaches are lower effort but have higher expressibility than rule- or template-based approaches, they tend to be less controllable to due the stochastic, black-box nature of large generative models. LLMs also often produce fairly repetitive, non-diverse outputs (Grunde-McLaughlin et al., 2023), and engineering prompts to guide LLMs toward desired outcomes can be a labor- and creativity-intensive process (Yeh et al., 2024; Zamfirescu-Pereira et al., 2023; Arawjo et al., 2024).
In our work, we aim to find a middle ground between structured and freeform text augmentation techniques (Figure 2). By applying a human-in-the-loop approach to data augmentation, we aim to help practitioners find relevant and diverse areas to augment through automatic, contextualized suggestions, but ultimately give them the agency to pick which directions to explore and which data to add. This reduces the effort required to design augmentation templates, while maintaining user controllability and expressibility.
@@ 2.2. Evaluating and Visualizing Text Diversity
We also take inspiration from general text visualization tools and techniques. There are several tools that visualize LLM text outputs to facilitate text comparison and sensemaking at scale (Arawjo et al., 2024; Kahng et al., 2024; Lilac, 2023; Gero et al., 2024). Specifically, clustering text embeddings is a common visualization approach to elucidate meaningful patterns in a dataset, e.g., topics or tasks (Siirtola et al., 2016; Pham et al., 2010; Smilkov et al., 2016). Building on these techniques, we contribute a human-in-the-loop approach to help ML practitioners both visualize and increase diversity through embedding visualizations. Additionally, our approach draws from works that design visual interfaces for prompt engineering tasks (e.g., (Feng et al., 2023; Almeda et al., 2024; Guo et al., 2024; Brade et al., 2023)) to help users produce diverse generative artificial intelligence (AI) outputs.
@@ 3. Formative Study
5 participants self-identified as ML researchers/engineers, 4 as research/engineering managers, and 3 as software engineers. Half of the practitioners worked on augmentation for text datasets (), while the remaining worked on other modalities including images, videos, and code. In addition, most participants augmented smaller datasets on the order of hundreds to a few thousands of data points (), in contrast to larger datasets with millions or billions of points (). When asked about this smaller data scale, F9 explained, “We often collect data targeted around a specific issue [that] requires finding very specific examples.” In terms of tasks, 6 participants worked on data augmentation for model fine-tuning (e.g., math reasoning and summarization), 4 worked on augmentation for model evaluation (e.g., risk assessment & mitigation, red teaming), and 2 worked on augmentation for model training.
@@ 3.1.1. Understand and explore topical data diversity (C1)
7 participants emphasized that a key challenge when augmenting data is “understanding the data and [its] problems,” especially in terms of data diversity (F12). F3 explained that a specific problem with unstructured text is that the “data has many dimensions to explore but it’s difficult to know which to look at.” Participants were interested in understanding topical diversity in particular, wanting to “capture things like topic modeling, breadth and representation” (F9). F1 shared that “we often need a detailed assessment [of topical diversity] that can’t be captured by existing metrics,” emphasizing the need for further exploration in this area. F6 also explained that when performing augmentation, “the number of data points isn’t as important, it’s more about the composition or representativeness of a dataset.” Overall, F2 expressed how “it would be most helpful to have a tool that allows engineers to quickly inspect and visualize data points themselves,” especially as it gets harder to assess diversity at scale.
@@ 3.1.2. Identify useful data instances and methods for augmentation (C2)
Related to C1, 5 participants noted that one of the most time-consuming and challenging parts of augmentation is determining which areas of the dataset would be most useful to augment. As F1 said, “How do you know what’s missing?” Another expert, F9, reported a similar sentiment for identifying effective augmentation approaches to use for model evaluation tasks: “A tool that offers new [and] less intuitive types of perturbations would be useful… and save a lot of time.” F3 elaborated on this thought, emphasizing that finding relevant augmentations can be difficult for text, where “semantic search is still lacking.” F7 agreed, as “with images, you can change their structure while maintaining semantics, but [that’s] really hard with text.” These findings echo previous work such as (Ding et al., 2024) and motivate the need for providing more actionable guidance to users during augmentation.
@@ 3.1.3. Augment data in an interpretable, controllable way (C3)
Another central theme was the need for more controllable and interpretable data augmentation techniques (). F8 told us that “interpretability is a huge aspect for product designers/engineers to act on data, and the lack of it feed[s] skepticism,” which is a struggle when using “more complex numerical approaches.” Similarly, F4 explained that many existing augmentation methods “were limited in context and confusing for our peers, who prefer more interpretable and simpler approaches.” With text, experts reported that manually writing examples is a common synthetic generation approach (). However, F2 noted that “human data collection is limited by time constraints so we can only do this on smaller scales.” Thus, there is a trend towards automated methods like LLMs due to the reduced human effort and costs (). A key concern with these latter approaches, however, is their lack of controllability and diversity. For example, “with LLMs, the structure of sentences tends to be repetitive and very cookie cutter… and the style of text [is] similar” (F12). The stochasticity of LLMs also makes their utility limited in many augmentation settings.
@@ 3.1.4. Ensure data quality while performing augmentation (C4)
Our experts also shared the challenges of ensuring that high quality data samples are added during augmentation (). As F8 pointed out, “It’s really about navigating the quality vs. scalability tradeoff… [adding data] might lower the quality, so you need to find the balance between collecting and curating data.” Participants F1 and F6 emphasized how difficult it is to evaluate the quality of generated data, as this usually requires extensive “manual inspection,” which is possible “with smaller datasets… [but] nearly impossible at larger scales.” F11 added that “it’s hard to have good metrics for measuring dataset quality. Like what’s the right mechanism to filter on, especially for things that are more heuristically driven rather than empirically,” which is often the case for text. In general, data quality requires a balance between diversity and relevance, as augmentation is often fairly targeted, e.g., to address a specific model deficit (F10), but also requires “making sure we’re covering our bases [and] achieving the right amount of diversity” (F5).
@@ 3.2.2. Detect low diversity and empty spaces in the data (T2)
Toward understanding data diversity (C1) and identifying useful areas for augmentation (C2), we also aim to help users find areas with low diversity and density. For example, participants mentioned wanting the ability to locate clusters with many “similar data points,” or those with only a few sentences (F2).
@@ 4. Augmentation Techniques
| Augmentation | Inputs | Automated Suggestions | Outputs |
@@
To support our goal of creating a more steerable and interpretable data augmentation process for unstructured text data, we develop three human-in-the-loop augmentation techniques: Augment with Concepts (Section 4.1), Augment by Interpolation (Section 4.2), and Augment with Large Language Model (LLM) (Section 4.3). We include different augmentation approaches to cater to various use cases and allow users to choose the method that best aligns with their specific needs and goals. Our augmentation techniques were crafted through an iterative design process, which started with 22 potential approaches, and each aims to address T3 by providing relevant augmentation suggestions to increase data diversity. We selected the final three methods based on perceived utility, novelty, and implementation feasibility.
An overview of all three approaches is shown in Figure 1 and summarized in Table 1. In this section, all example sentences are drawn from a dataset containing CHI 2024 paper titles (Guerra, 2024) and are illustrated in our system interface in Section 5. To begin, each augmentation strategy takes as input a single sentence , along with its corresponding embedding and the number of desired new sentences , to generate output sentences .
@@
Our first two methods — Augment with Concepts (Section 4.1) and Augment by Interpolation (Section 4.2) — are embedding-based augmentation approaches. Both methods leverage Vec2Text (Morris et al., 2023), an embedding inversion technique based on controlled text generation (Hu et al., 2017). In short, an embedding inversion converts a high-dimensional vector back into text.
@@ 4.1. Augment with Concepts
New Sentence: Cardinal Cards: An Engaging Card-Based Method for AI-Driven Statistical Data Exploration
@@ 4.2. Augment by Interpolation
In our interpolation approach, the user must select a second sentence to help augment . The user can take one of Amplio’s suggestions for (selecting ), or choose their own sentence (see Section 5.2.2). Drawing from existing text embedding interpolation techniques (e.g., (Chen et al., 2020; Seyler and Zhai, 2020; Kilbas et al., 2024)), we perform linear interpolation between a start embedding and an end embedding , which correspond to the two input sentences ( and , respectively). We can then generate output embeddings by creating equally spaced values (i.e., ), which control the weight scalars used for interpolation. Thus, each output is computed as:

### v2

@@ 1. Introduction
In machine learning (ML), data plays a key role in driving model behavior. Even the most sophisticated, specialized model architectures may underperform when training data is limited or if data instances are low quality and noisy (Xie et al., 2020; Shorten and Khoshgoftaar, 2019; Sambasivan et al., 2021; Lara and Tiwari, 2022). As such, data augmentation — the process of creating new data samples to add to an existing dataset — is a common and important practice for numerous ML applications (e.g., model training (Rebuffi et al., 2021; Khosla and Saini, 2020) and fine-tuning (Feng et al., 2020; Shi et al., 2022)). In particular, data augmentation plays a critical role in generating novel use cases for model evaluation, enabling tasks such as perturbation analyses (Shorten and Khoshgoftaar, 2019; Xie et al., 2020; Sui et al., 2024), fairness testing (Navarro et al., 2024; Zhang and Sang, 2020), and red teaming attacks (Feffer et al., 2024; Ganguli et al., 2022). Across many of these scenarios, a shared goal is to improve dataset diversity to make ML models more robust and safe (Chao et al., 2023; Shi et al., 2022; Rebuffi et al., 2021), which is especially important given the growing complexity and applications of these models.
@@
The ease and efficacy of data augmentation is also highly dependent on data modality. For instance, with structured data like tables, it is relatively straightforward to determine where augmentation might be needed by looking at the distribution of features (i.e., categories) and adding new rows. Tabular data can also be augmented by adding entirely new features to the table (i.e., columns). On the other hand, it is much harder to augment unstructured data, like images or text (Figure 1). Such modalities do not have the same kind of identifiable “features” to augment along. Furthermore, for text data, it is non-obvious how to perform “distortions” to natural language without altering its semantics or context (Ding et al., 2024; Qu et al., 2021), whereas for images, it is easier to apply rotation, cropping, and brightness modifications to augment data. With unstructured text, it is also less clear what kinds of diversity to strive for when performing augmentation, as there are many axes to consider, including topical, syntactic, and lexical diversity (Chen et al., 2020; Reif et al., 2023).
@@
With these augmentation methods, we designed and developed Amplio, an interactive data augmentation tool for unstructured text datasets. Amplio is designed to help ML practitioners systematically navigate “unknown unknowns” and diversify their datasets by finding relevant empty data spaces (i.e., parts of the desired dataset distribution with few or no data points) to explore. Our tool visualizes sentences in an embedding plot, where literal empty regions serve as a metaphor for under-explored areas in the data distribution, and works to support augmentation processes by providing an interface for users to fill in these data gaps using our three techniques (Figure 1). To assess these augmentation approaches, we conducted a user study designed around red teaming LLMs, a common and important real-world data augmentation task for model evaluation, as identified by our formative study. We recruited 18 professional red teamers to augment a harmful LLM prompts dataset (Bai et al., 2022) using Amplio, finding that our augmentation methods and visualizations were effective in generating relevant, diverse data. Participants also discovered unique use cases for each augmentation technique, suggesting additional design opportunities for interactive data augmentation tools.
@@
A suite of three human-in-the-loop data augmentation techniques—Augment with Concepts, Augment by Interpolation, and Augment with LLM—to help users find diverse and relevant “empty spaces” to explore when augmenting unstructured text data.
The design and implementation of Amplio, an interactive visualization tool that applies our three techniques to help users augment text datasets in a controllable and interpretable way.
@@ 2. Related Work
Visualization plays a key role in understanding and evaluating ML models (Brath et al., 2023; Beauxis-Aussalet et al., 2021). Many existing tools focus on exploring data distributions and quality (Research, 2021a, 2017; Face, 2021; IBM, 2021; Assogba et al., 2023; Giesen et al., 2017), understanding data iteration (Hohman et al., 2020), or assessing factors like model fairness (Ahn and Lin, 2019; Bird et al., 2020; Xie et al., 2020; Wexler et al., 2019) and interpretability (Hohman et al., 2019; Tenney et al., 2020; Wang et al., 2020b; Strobelt et al., 2018). There is also interest in designing visualization techniques to aid developers in creating safe, trustworthy, and responsible models (Reif et al., 2023; Templeton, 2024; Beauxis-Aussalet et al., 2021; Feng et al., 2024; Wang et al., 2024).
In contrast, our work targets data augmentation processes, using human-in-the-loop methods to improve data diversity for unstructured text. Specifically, we focus on synthetic text augmentation.
@@ 2.1. Synthetic Text Data Augmentation
Recently, language models have emerged as an alternative mechanism for data space augmentations (Guo and Chen, 2024; Wagner et al., 2024; Patel et al., 2024; Li et al., 2023; Peng et al., 2023). For example, LLMs have been used to generate datasets of counterfactual sentences (Wu et al., 2021) and adversarial prompts (Samvelyan et al., 2024). While these approaches are lower effort and have higher expressibility than rule- or template-based approaches, they tend to be less controllable to due the stochastic, black-box nature of large generative models. LLMs also often produce fairly repetitive, non-diverse outputs (Grunde-McLaughlin et al., 2023), and engineering prompts to guide these models toward desired outcomes can be a labor- and creativity-intensive process (Yeh et al., 2024; Zamfirescu-Pereira et al., 2023; Arawjo et al., 2024).
Slider showing how our system, Amplio, aims to provide a middle ground between freeform and structured text augmentation approaches.
In our work, we strive to find a middle ground between structured and freeform text augmentation techniques (Figure 2). By applying a human-in-the-loop approach, we aim to help practitioners find relevant and diverse areas of datasets to augment through automatic, contextualized suggestions. However, ultimately, practitioners retain the agency to pick which directions to explore and which data to add. This reduces the effort required to design augmentation templates and prompts, while maintaining user controllability and expressibility.
@@ 2.2. Evaluating and Visualizing Text Diversity
We also take inspiration from general text visualization tools and techniques. There are several tools that visualize LLM text outputs to facilitate text comparison and sensemaking at scale (Arawjo et al., 2024; Kahng et al., 2024; Lilac, 2023; Gero et al., 2024). Specifically, clustering text embeddings is a common visualization approach to summarize and elucidate meaningful patterns in a dataset, e.g., topics or tasks (Siirtola et al., 2016; Pham et al., 2010; Smilkov et al., 2016). Building on these techniques, we contribute a human-in-the-loop approach to help ML practitioners both visualize and increase diversity through embedding visualizations. Additionally, our approach draws from works that design visual interfaces for prompt engineering tasks (e.g., (Feng et al., 2023; Almeda et al., 2024; Guo et al., 2024; Brade et al., 2023)) to help users produce diverse generative artificial intelligence (AI) outputs.
@@ 3. Formative Study
Five participants self-identified as ML researchers/engineers, 4 as research/engineering managers, and 3 as software engineers. Half of the practitioners worked on augmentation for text datasets (), while the remaining worked on other modalities including images, videos, and code. In addition, most participants augmented smaller datasets on the order of hundreds to a few thousands of data points (), in contrast to larger datasets with millions or billions of points (). When asked about this smaller data scale, F9 explained, “We often collect data targeted around a specific issue [that] requires finding very specific examples.” In terms of tasks, 6 participants worked on data augmentation for model fine-tuning (e.g., math reasoning and summarization), 4 worked on augmentation for model evaluation (e.g., risk assessment & mitigation, red teaming), and 2 worked on augmentation for model training.
@@ 3.1.1. Understand and explore topical data diversity (C1)
Seven participants emphasized that a key challenge when augmenting data is “understanding the data and [its] problems,” especially in terms of data diversity (F12). F3 explained that a specific problem with unstructured text is that the “data has many dimensions to explore but it’s difficult to know which to look at.” Participants were most interested in understanding topical diversity, wanting to “capture things like topic modeling, breadth and representation” (F9). F1 shared that “we often need a detailed assessment [of topical diversity] that can’t be captured by existing metrics,” emphasizing the need for further exploration in this area. F6 also explained that when performing augmentation, “the number of data points isn’t as important, it’s more about the composition or representativeness of a dataset.” Overall, F2 expressed how “it would be helpful to have a tool that allows engineers to quickly inspect and visualize data points themselves,” especially as it gets harder to assess diversity at scale.
@@ 3.1.2. Identify useful data instances and methods for augmentation (C2)
Related to C1, five participants noted that one of the most time-consuming and challenging parts of augmentation is determining which areas of the dataset would be most useful to augment. As F1 said, “How do you know what’s missing?” Another expert, F9, reported a similar sentiment for identifying effective augmentation approaches to use for model evaluation tasks: “A tool that offers new [and] less intuitive types of perturbations would be useful… and save a lot of time.” F3 elaborated on this thought, emphasizing that finding relevant augmentations can be difficult for text, where “semantic search is still lacking.” F7 agreed, as “with images, you can change their structure while maintaining semantics, but [that’s] really hard with text.” These findings echo previous work such as (Ding et al., 2024) and motivate the need for providing more actionable guidance to users during augmentation.
@@ 3.1.3. Augment data in an interpretable, controllable way (C3)
Another central theme was the need for more controllable and interpretable data augmentation techniques (). F8 told us that “interpretability is a huge aspect for product designers/engineers to act on data, and the lack of it feed[s] skepticism,” which is a struggle when using “more complex numerical approaches.” Similarly, F4 explained that many existing augmentation methods “were limited in context and confusing for our peers, who prefer more interpretable and simpler approaches.” With text, experts reported that manually writing examples is a common synthetic generation approach (). However, F2 noted that “human data collection is limited by time constraints so we can only do this on smaller scales.” Thus, there is a trend towards automated methods like LLMs due to the reduced human effort and costs (). A key concern with these latter approaches, however, is their lack of controllability and diversity. For example, “with LLMs, the structure of sentences tends to be repetitive and very cookie cutter… and the style of text [is] similar” (F12). The stochasticity of LLMs also makes their utility limited in many augmentation settings (Zamfirescu-Pereira et al., 2023; Yeh et al., 2024).
@@ 3.1.4. Ensure data quality while performing augmentation (C4)
Our experts also shared the challenges of ensuring that high quality data samples are added during augmentation (). As F8 pointed out, “It’s really about navigating the quality vs. scalability tradeoff… [adding data] might lower the quality, so you need to find the balance between collecting and curating data.” Participants F1 and F6 emphasized how difficult it is to evaluate the quality of generated data, as this usually requires extensive “manual inspection,” which is feasible “with smaller datasets… [but] nearly impossible at larger scales.” F11 added that “it’s hard to have good metrics for measuring dataset quality. Like what’s the right mechanism to filter on, especially for things that are more heuristically driven rather than empirically,” which is often the case for text. In general, participants expressed that data quality requires a balance between diversity and relevance, as augmentation is often fairly targeted, e.g., to address a specific model deficit (F10), but also requires “making sure we’re covering our bases [and] achieving the right amount of diversity” (F5).
@@ 3.2.2. Detect low diversity and empty spaces in the data (T2)
Toward understanding data diversity (C1) and identifying useful areas for augmentation (C2), we also aim to help users find areas with low diversity and density. For example, participants mentioned wanting the ability to locate clusters with many “similar data points,” or those with only a few sentences (F2). We use the latter to shape our definition of “empty spaces” in this work.
@@ 4. Augmentation Techniques
| Augmentation by | Inputs | Automated Suggestions | Outputs |
@@
Summary of inputs, automated suggestions, and outputs of our proposed human-in-the-loop text augmentation methods.
To support our goal of creating a more steerable and interpretable data augmentation process for unstructured text data, we develop three human-in-the-loop augmentation techniques: Augment with Concepts (Section 4.1), Augment by Interpolation (Section 4.2), and Augment with Large Language Model (LLM) (Section 4.3). We include different augmentation approaches to cater to various use cases and allow users to choose the method that best aligns with their specific needs and goals. Our augmentation techniques were crafted through an iterative design process with 22 initial approaches that varied by input/output configuration (e.g., one-to-one vs. one-to-many), interaction modality (e.g., drawing vs. natural language inputs), and degree of automation (Figure 2). We selected our final three methods based on perceived utility, novelty, generation speed, and implementation feasibility. Each technique aims to address T3 by providing relevant augmentation suggestions to increase data diversity.
Dataset: CHI 2024 Paper Titles
An overview of all three approaches is shown in Figure 1 and summarized in Table 1. In this section, all example sentences are drawn from a dataset containing CHI 2024 paper titles (Guerra, 2024) and are illustrated in our system interface in Section 5. To begin, each augmentation strategy takes as input a single sentence , along with its corresponding embedding and the number of desired new sentences , to generate output sentences . Throughout the paper, we box example data points and their augmentations to demonstrate our techniques.
@@
Our first two methods — Augment with Concepts (Section 4.1) and Augment by Interpolation (Section 4.2) — are embedding-based augmentation approaches. Both methods leverage Vec2Text (Morris et al., 2023), a state-of-the-art embedding inversion technique based on controlled text generation (Hu et al., 2017). In short, an embedding inversion converts a high-dimensional vector back into text.
@@ 4.1. Augment with Concepts
→ New Sentence: Cardinal Cards: An Engaging Card-Based Method for AI-Driven Statistical Data Exploration
@@ 4.2. Augment by Interpolation
In our interpolation approach, the user selects a second sentence to help augment . The user can take one of Amplio’s suggestions for (selecting ), or choose their own sentence (see Section 5.2.2). Drawing from existing text embedding interpolation techniques (e.g., (Chen et al., 2020; Seyler and Zhai, 2020; Kilbas et al., 2024)), we perform linear interpolation between a start embedding and an end embedding , which correspond to the two input sentences ( and , respectively). We can then generate output embeddings by creating equally spaced values (i.e., ), which control the weight scalars used for interpolation. Thus, each output is computed as: